OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). tan x + sec x = 1

OpenStudy (anonymous):

x = 0 x = No solution x =

OpenStudy (jhannybean):

@dan815

OpenStudy (anonymous):

sec(x) + tan(x) = 1 1/cos(x) + sin(x)/cos(x) = 1 (1 + sin(x)) / cos(x) = 1 1 + sin(x) = cos(x) 1 = cos(x) - sin(x) x = 0 radians, 2π, 4π ... 2nπ radians But since your interval only allows for the answer 0, then that's your only solution.

OpenStudy (anonymous):

wait is it 0 π/4 5π/4 or no solution

OpenStudy (anonymous):

 They share commonalities at 0 radians, so there you go!